A converging bundle of light rays in the shape of cone with a vertex angle of 45 degrees falls on a circular diaphragm of 20cm diametre.A lens with power 5D is fixed in the diaphragm .diameter of face of lens is equal to that of diaphragm. If the vertex angle of new cone is theta then tan theta = 

Grade:9

2 Answers

Gundeep Singh
7 years ago
I think it is 4/3 but i m not sure. I have assumed that the rays are comming from infinity so they converge at focus i.e at a distance of 20cm from lens. Radius of lens =10cm. So tan(theeta/2) =10/20 =1/2 And tan (theeta) =4/3.
suprabha
5 years ago
A converging bundle of light rays in the shape of cone with a vertex angle of 45 degrees falls on a circular diaphragm of 20cm diametre.A lens with power 5D is fixed in the diaphragm .diameter of face of lens is equal to that of diaphragm. If the vertex angle of new cone is theta then tan theta =